### v1

# Out-of-Distribution Semantic Occupancy Prediction

Abstract
3D Semantic Occupancy Prediction is crucial for autonomous driving, providing a dense, semantically rich environmental representation. However, existing methods focus on in-distribution scenes, making them susceptible to Out-of-Distribution (OoD) objects and long-tail distributions, which increases the risk of undetected anomalies and misinterpretations, posing safety hazards. To address these challenges, we introduce Out-of-Distribution Semantic Occupancy Prediction, targeting OoD detection in 3D voxel space. To fill the gaps in the dataset, we propose a Synthetic Anomaly Integration Pipeline that injects synthetic anomalies while preserving realistic spatial and occlusion patterns, enabling the creation of two datasets: VAA-KITTI and VAA-KITTI-360. We introduce OccOoD, a novel framework integrating OoD detection into 3D semantic occupancy prediction, with Voxel-BEV Progressive Fusion (VBPF) leveraging an RWKV-based branch to enhance OoD detection via geometry-semantic fusion. Experimental results demonstrate that OccOoD achieves state-of-the-art OoD detection with an AuROC of 67.34% and an AuPRCr of 29.21% within a 1.2m region, while maintaining competitive occupancy prediction performance. The established datasets and source code will be made publicly available at https://github.com/7uHeng/OccOoD.
1 Introduction
3D scene understanding is fundamental to autonomous driving, crucial for motion planning and obstacle avoidance [36]. The cost-effectiveness and ease of deployment of cameras have made vision-based 3D understanding methods a focal point of research [12, 66, 121, 37, 114, 97, 82, 53, 62]. Among these methods, 3D semantic occupancy prediction networks [41, 51, 66] are distinguished by their ability to accurately reconstruct geometry and semantic details. By producing detailed occupancy maps with rich semantic context, these networks improve environmental perception and decision-making, enhancing the safety and reliability of autonomous navigation.
However, current methods [41, 85, 51] primarily focus on in-distribution scenarios, where models are trained and evaluated on predefined object categories and environments. While effective in controlled settings, this approach [53, 12] exposes systems to vulnerabilities from OoD objects and anomalies commonly encountered in real-world scenarios [55, 8, 33, 17], as seen in Fig. LABEL:fig:teaser where the standard semantic occupancy prediction method [66] produces misclassifications. OoD detection—the capability to identify entities that substantially deviate from the training data distribution—remains a critical yet underexplored challenge. Such anomalies include unexpected obstacles (e.g., discarded furniture, stray wildlife, construction equipment), unusual road conditions (e.g., potholes, debris, or oil spills), and atypical weather phenomena (e.g., sudden fog, heavy rain, or snowdrifts). As existing systems may misinterpret or fail to detect these deviations, they pose significant safety risks, potentially leading to failures and even catastrophic consequences in autonomous driving.
To bridge this critical gap, we introduce the task of Out-of-Distribution Semantic Occupancy Prediction. However, a key challenge in this domain is the scarcity of datasets annotated for both semantic occupancy and OoD anomalies, hindering robust model development. To address this, we propose a Synthetic Anomaly Integration Pipeline that injects diverse anomalies into existing datasets while preserving realistic spatial and occlusion patterns. This pipeline enables the creation of two novel datasets comprising $26$ distinct categories of anomalies, where anomaly annotations encompass a diverse range of unforeseen hazards, facilitating comprehensive evaluation of OoD detection in 3D scene understanding. Building upon our newly proposed datasets for out-of-distribution detection in semantic occupancy prediction, we introduce OccOoD, the first unified framework that seamlessly integrates OoD detection into 3D semantic occupancy prediction.
OccOoD incorporates several key innovations to enhance anomaly detection in complex environments. First, existing methods [66, 51] rely on a single-voxel representation, which limits the effectiveness of OoD detection. To address this, we propose a Voxel-BEV Progressive Fusion strategy, which integrates voxel-based and Bird’s-Eye-View (BEV) representations. BEV enhances global scene understanding, enabling improved occupancy prediction and robust OoD detection. Second, we introduce a dedicated OoD detection mechanism leveraging entropy and cosine similarity, enabling effective identification of anomalies in dynamic and unstructured scenes. Finally, to ensure comprehensive evaluation, we conduct extensive OoD detection experiments on the newly proposed datasets and verify our occupancy prediction performance based on public benchmarks. The experimental results show that OccOoD achieves AuROC scores of $65.91\%$ and $67.34\%$ on the two proposed datasets, respectively, and attains mIoU scores of $15.85\%$ and $18.37\%$ on SemanticKITTI [3], and KITTI-360 [52], respectively. These results collectively establish OccOoD as a robust and scalable solution for addressing OoD challenges in 3D semantic occupancy prediction.
In a nutshell, our contributions are three-fold:
  • We for the first time introduce out-of-distribution detection into the task of semantic occupancy prediction by delivering two new datasets VAA-KITTI and VAA-KITTI-360, specifically designed to evaluate out-of-distribution detection in the context of autonomous driving safety.
  • We propose OccOoD, a multi-view fusion framework that integrates front-view and bird’s-eye-view representations, enabling a comprehensive spatial understanding of outliers and seamless integration of OoD detection into 3D semantic occupancy prediction. This framework robustly handles both known and unknown objects.
  • Our proposed framework, OccOoD, achieves state-of-the-art OoD detection performance on the established datasets while maintaining competitive accuracy in 3D semantic occupancy prediction.
2 Related Work
Semantic Occupancy Prediction. 3D Semantic Occupancy Prediction aims to infer both the geometry and semantics of a scene by inferring the occupancy state of voxels [100, 73, 19, 79, 105]. To advance estimation quality and efficiency, efforts are made in 3D representation [23, 56, 62, 35, 78, 98, 82, 117, 39, 49, 61, 44, 42], context enhancement [101, 86, 46, 89, 124, 96, 48], modality fusion [30, 113, 70, 14, 75, 64, 104, 120], temporal association [92, 69, 91, 109, 108, 11], long-sequence and generative modeling [54, 90, 123, 47, 85], self-supervision [38, 32, 10, 9, 71], and offboard processing [63, 77, 94, 118, 40]. These efforts have collectively improved the quality and efficiency of semantic occupancy prediction. Recently, several works [121, 41, 66, 110] have introduced more fine-grained feature processing techniques. Meanwhile, uncertainty-aware methods [93, 80, 13] incorporate uncertainty quantification to enhance estimation reliability. However, these methods are limited to predefined categories, requiring retraining to generalize to novel object types and lacking flexibility. While open-vocabulary learning [81, 119, 9, 87, 116] addresses generalization to arbitrary categories, it remains constrained to In-Distribution (ID) scenarios and fails to handle OoD detection. In real-world applications, the categories of anomalous objects are unforeseeable, making such methods less reliable. In contrast, our proposed model extends semantic occupancy prediction by perceiving all OoD objects, enabling efficient and comprehensive OoD detection.
Figure 1: An illustration of the proposed Synthetic Anomaly Integration Pipeline. This pipeline is designed to address the challenges of collecting real-world OoD data by synthesizing anomalies that are both physically plausible and contextually realistic.
Out-of-Distribution Segmentation. Out-of-distribution segmentation identifies anomalies in specific scenes [22, 17, 55, 8] while maintaining in-distribution segmentation capabilities. Current research primarily focuses on anomaly detection in 2D images of road scenes, with the main methods including per-pixel architectures and mask-transformer-based approaches [1, 29, 74, 68, 25]. The per-pixel architecture methods include uncertainty-based [67, 21, 33, 43, 16], reconstruction-based [55, 99, 31, 88, 27], and outlier exposure [18, 84, 28, 58, 20] methods. Although these methods have advanced 2D anomaly detection, they are limited to narrow field-of-view pinhole cameras, hindering their application to complex 3D scenarios requiring high-precision spatial perception. Some studies extend anomaly detection to 3D space; however, these efforts primarily focus on industrial [5, 57] or indoor object defect detection [7, 50], where models fuse static point cloud data to optimize defect localization [34, 4, 15, 122, 111]. Due to reliance on geometric consistency assumption of static objects and limited sensitivity to dynamic changes, these methods fail to meet the demands of more complex scenarios. In comparison, our work introduces OoD segmentation into vision-based 3D semantic occupancy prediction, overcoming the limitations of current methods in dynamic environments.
3 Out-of-Distribution Datasets
We introduce the task of Out-of-Distribution Semantic Occupancy Prediction. To address the lack of evaluation platforms, we propose a Synthetic Anomaly Integration Pipeline to generate abnormal objects in real road scene images, leading to the creation of two novel datasets: VAA-KITTI and VAA-KITTI-360.
3.1 Synthetic Anomaly Integration Pipeline
Collecting images and point cloud data of OoD objects is challenging due to their rarity in real-world scenes and the high cost of data collection and annotation. To address this, we propose a Synthetic Anomaly Integration Pipeline that generates synthetic anomalies under physical and environmental constraints, ensuring plausibility and challenge for robust OoD detection model evaluation. As illustrated in Fig. 1, our pipeline is structured into three key phases:
Phase 1: Anomaly Generation and Localization. In 2D anomaly detection, studies have demonstrated that synthetic anomaly data is highly effective for model training and evaluation [24, 59]. We employ POC [24], a state-of-the-art image generation technique, to synthesize anomaly images and their corresponding mask maps, enabling precise localization of anomalous objects. To better align with real-world application scenarios, we introduce multi-category anomalies into the images, as shown in Fig. 2.
Phase 2: Depth Estimation and 3D Transformation. This phase focuses on integrating image data with accurate depth information, facilitating the transformation of 2D anomaly detection results into precise 3D spatial localization. To address the challenge of limited metric depth estimation accuracy, we propose a method to project synthesized anomalies into the 3D point cloud.
We first utilize the large version of Depth Anything V2 [103], a state-of-the-art depth estimation model, to compute pseudo-depth for both original and synthesized images. Subsequently, we randomly sample a predefined number of points, from the point cloud and project them onto the image plane to obtain pixel coordinates $\mathbf{u}_{i}$ and Euclidean distances $d_{i}$ as:
$\begin{split}\mathbf{u}_{i}&=\frac{\mathbf{K}\cdot(\mathbf{R}\mathbf{p}_{i}+\mathbf{t})}{[\mathbf{R}\mathbf{p}_{i}+\mathbf{t}]_{z}},\\ d_{i}&=\|\mathbf{R}\mathbf{p}_{i}+\mathbf{t}\|_{2},\end{split}$ (1)
where $\mathbf{p}_{i}$ is a 3D point, $\mathbf{K}$ is the camera intrinsic matrix, $\mathbf{R}$ and $\mathbf{t}$ are the rotation and translation of the camera, $\mathbf{u}_{i}$ is the projected 2D point, and $d_{i}$ is the Euclidean distance.
By utilizing $\mathbf{u}_{i}$ and the previously obtained depth map, we align the pseudo-depth with the real depth and apply Support Vector Regression (SVR) [2] for nonlinear regression, mapping the pseudo-depth to real-world depth. Finally, the aligned depth map and mask map from Phase 1 are used to derive 3D anomaly coordinates $(u,v,d)$, which are then transformed into the 3D voxel space.
Phase 3: Voxelization and Scene Integration. The pseudo-point clouds of anomalies are voxelized and integrated into the 3D voxel space of the original point cloud. Leveraging camera poses, we project each anomaly point $\mathbf{A}_{i}^{v}=(x_{i}^{v},y_{i}^{v},z_{i}^{v})$ (representing the discrete voxel coordinates of the $i$-th anomaly) back into the scene to determine occlusion relationships via ray marching. The ray position of the ray at step $k$ is computed as:
$\mathbf{c}_{k}=\mathbf{C}^{v}+k\cdot\frac{\mathbf{A}_{i}^{v}-\mathbf{C}^{v}}{\|\mathbf{A}_{i}^{v}-\mathbf{C}^{v}\|_{2}},$ (2)
where $k$ is the step index, $\mathbf{c}_{k}$ is the ray position at $k$, $\mathbf{C}^{v}$ is the camera position and $\mathbf{v}_{k}=\text{round}(\mathbf{c}_{k})$ is the corresponding voxel center. The step size is controlled by a scaling factor $s>0$, with $\Delta=\frac{1}{s}$ defining the resolution.
The occlusion process is formalized as:
$\mathbf{V}[\mathbf{v}_{k}]=\begin{cases}1,&\text{if }\|\mathbf{c}_{k}-\mathbf{A}_{i}^{v}\|_{2}\leq\Delta,\\ 0,&\text{otherwise}.\end{cases}$ (3)
where 00 and $1$ represent the occupancy status of the voxel.
3.2 VAA-KITTI and VAA-KITTI-360
Figure 2: Distribution map of VAA-KITTI and VAA-KITTI-360 datasets. The datasets contain $26$ distinct OoD categories, which can be divided into five main categories as shown in the figure.
We applied the aforementioned synthesis pipeline to the SemanticKITTI [3] and SSCBench-KITTI-360 [52] datasets, resulting in the creation of two new synthetic datasets: VAA-KITTI and VAA-KITTI-360. These synthesized datasets extend the original label sets by introducing anomaly labels, encompassing $26$ distinct categories of anomalies such as animals, furniture, and garbage bags.
To ensure the plausibility and naturalness of the generated images, over $30,000$ images produced using the POC algorithm were carefully reviewed by three independent annotators. After multiple rounds of cross-validation, images with poor generation quality, implausible object placements, or non-anomalous scenarios (e.g., birds flying in an urban sky or trash, trash bags, and wheeled bins placed alongside sidewalks) are excluded, resulting in a final selection of $500$ images for both the VAA-KITTI and VAA-KITTI-360 datasets to form the test set for the Out-of-Distribution Semantic Occupancy Prediction task.
Regional-Level Metrics Design: Referring to the field of 2D image anomaly detection, we adopt Area under the Receiver Operating Characteristic Curve (AuROC) as one of the evaluation metrics for model performance. Furthermore, to mitigate the evaluation impact caused by the inaccuracy of existing methods in predicting the absolute positions of instances, and bridge the gap between localization demands and anomaly-awareness goals, we propose a regional-level evaluation metric: Regional Area under the Precision-Recall Curve ($\text{AuPRC}_{r}$). The $\text{AuPRC}_{r}$ incorporates tolerance for minor localization inaccuracies while maintaining sensitivity to anomaly detection.
$\text{AuPRC}_{r}$: The Precision-Recall (PR) curve evaluates the balance between precision and recall at varying thresholds:
$\text{Precision}_{r}=\frac{\text{TP}_{r}}{\text{TP}_{r}+\text{FP}_{r}},\quad\text{Recall}_{r}=\frac{\text{TP}_{r}}{\text{TP}_{r}+\text{FN}_{r}},$ (4)
where $\text{TP}_{r}$, $\text{FP}_{r}$, and $\text{FN}_{r}$ denote true positives, false positives, and false negatives within a regional tolerance radius $r$. The $\text{AuPRC}_{r}$ is computed as Eq. (5):
$\text{AuPRC}_{r}=\int_{0}^{1}\text{Precision}_{r}(\text{Recall}_{r})\,d(\text{Recall}_{r}).$ (5)
4 OccOoD: Proposed Framework
4.1 Problem Setting
Out-of-distribution segmentation in 2D pinhole images typically relies on pixel-wise classification scores. Building on this concept, we extend out-of-distribution detection to 3D semantic occupancy prediction at the voxel level. Given an input image $I\in\mathbb{R}^{C\times H\times W}$, the semantic occupancy prediction process yields a volumetric representation $V\in\mathbb{R}^{C\times H\times W\times Z}$, where $Z$ denotes the depth dimension. The voxel-level classification scores $Q(V)\in\mathbb{R}^{K\times H\times W\times Z}$ are then obtained. To assess anomalies at the voxel level, an anomaly score $A(V)\in\mathbb{R}^{H\times W\times Z}$ is computed as Eq. (6):
$A(V)=f(Q(V)),$ (6)
where $K$ denotes the number of classes for the semantic prediction, $C$, $H$, $W$, and $Z$ correspond to the channels, height, width, and depth of the voxel space, respectively.
4.2 Overall Architecture
Figure 3: An overview of the proposed OccOoD framework. The image encoder extracts 2D features and converts them into 3D features via View Transformation, guided by geometric occupancy. Seed voxels and other features are processed through geometry pairing and semantic refinement to generate enhanced BEV and voxel features. Finally, cross-view feature synergy integrates these representations for rich geometry and high spatial efficiency.
The OccOoD framework, illustrated in Fig. 3, integrates front-view and BEV representations for precise 3D semantic occupancy prediction. While 3D voxels preserve fine-grained geometric details, BEV enhances spatial reasoning and global scene understanding. To effectively merge these representations, we propose Voxel-BEV Progressive Fusion (VBPF), which iteratively refines voxel-BEV integration, leveraging voxel geometry and BEV efficiency for improved feature representation. This multi-perspective fusion enhances generalization and robustness, significantly boosting out-of-distribution detection performance. Initially, following SGN [66], an image encoder extracts 2D features from RGB images, which are then mapped into 3D space using a view transformation module adapted from MonoScene [12]. The coarse 3D features are refined through geometry guidance, and an auxiliary 3D occupancy head is optimized using binary cross-entropy loss:
$\small L_{\text{geo}}=-\sum_{i}\left[(1-Y_{o,i})\log(1-\hat{Y}_{o,i})+Y_{o,i}\log(\hat{Y}_{o,i})\right],$ (7)
where $i$ indexes voxels, $Y_{o}$ represents ground truth occupancy, and $\hat{Y}_{o}$ indicates the predicted occupancy.
After the proposal network filters the 3D features, optimized with a binary cross-entropy loss $L_{\text{occ}}$, the filtered features undergo Geometry Parsing and Semantic Refinement to extract multi-scale BEV features and fine-grained semantic voxel features. These BEV features are fused with the original 3D features through a UNet-like module, ensuring precise alignment and enhanced representation learning. Finally, the BEV representation is upsampled, concatenated with the aggregated voxel representation, and refined using convolutions to produce the semantic occupancy prediction.
4.3 Voxel-BEV Progressive Fusion
The coarse voxel features $F^{3D}$ from the view transformation module and the seed features $F_{\text{s,0}}^{3D}$ obtained through the proposal network are both 3D voxel features. However, a single representation limits fine-grained semantics and segmentation boundaries, hindering OoD detection. To overcome this, VBPF progressively integrates voxel-based and BEV representations, leveraging BEV’s top-down perspective for global context and spatial uncertainty identification.
Geometry Parsing: The coarse voxel features $F^{3D}$ are processed through a geometric parsing branch, comprising a linear layer, residual connections, a BEV projection module, and a Vision-RWKV (VRWKV) module [26] (see Fig. 3), to encode rich geometric information into compact BEV representations.
Specifically, $F^{3D}$ is initially projected into a lower-dimensional space via a linear layer, refined through a residual block, aligned along the z-axis, and compressed using 2D convolutions to generate dense BEV features. These BEV features are further refined by the VRWKV module, which enhances their representational quality through spatial and channel mixing operations, producing dense geometric BEV features $\{F_{\text{g}_{i}}^{\text{BEV}}\}_{i}$.
Semantic Refinement: The seed voxels $F_{\text{s,0}}^{3D}$ generated by the proposal network undergo semantic refinement to enhance semantic accuracy. The semantic refinement module comprises three cascaded sparse encoder blocks (SEB), each linked with a BEV projection module and a VRWKV block. Each SEB contains a Sparse Feature Encoder (SFE) [107] and a Sparse Geometry Feature Encoder (SGFE) [107], which optimize the initial semantic representation and refine it through hierarchical pooling and attention mechanisms, capturing fine-grained local details and broader contextual dependencies.
First, the voxel features pass through the SEB, where two SEBs output contextual features at different scales, $F_{\text{s,1}}^{3D}$ and $F_{\text{s,2}}^{3D}$, each embedding rich geometric and contextual information. These features are projected into BEV space via a BEV projection operation, where max-pooling aggregates features with the same BEV index, generating sparse BEV features, which are then converted into dense semantic-BEV features using Spconv’s densification function.
The VRWKV module is then applied to model long-range dependencies, enhancing the representational quality of the BEV features. The final output includes multi-scale semantic BEV features: $\{F_{\text{s}_{i}}^{\text{BEV}}\}_{i}$, as well as the refined semantic feature $F_{\text{s}}^{3D}$ obtained by Eq. (8):
$F_{\text{s}}^{3D}=\text{MLP}([F_{\text{s,0}}^{3D},F_{\text{s,1}}^{3D},F_{\text{s,2}}^{3D}]).$ (8)
The multi-scale semantic BEV features $\{F_{\text{s}_{i}}^{\text{BEV}}\}_{i}$ are used for BEV-level fusion, while the refined semantic feature $F_{\text{s}}^{3D}$ is utilized for voxel-level fusion. Additionally, $F_{\text{s}}^{3D}$ is passed through an auxiliary semantic head (a two-layer MLP) to predict the corresponding semantics $\hat{Y}_{s}$.
The semantic refinement stage is optimized using a combination of cross-entropy loss and Lovasz loss [6]:
$L_{sem}=L_{ce}(\hat{Y}_{s},Y_{s})+L_{lovasz}(\hat{Y}_{s},Y_{s}),$ (9)
where $Y_{s}$ represents the semantic labels of seed voxels indexed from the semantic scene labels $Y$.
Cross-View Feature Synergy: To effectively integrate BEV and voxel-level representations, we first extract features independently. For BEV processing, we employ a UNet-like architecture, where the encoder captures hierarchical spatial features and the decoder reconstructs high-resolution feature maps for seamless fusion. The inputs include the projection of 3D features into BEV space ($F^{\mathrm{BEV}}_{c}$), geometric BEV features $\{F_{\text{g}_{i}}^{\text{BEV}}\}_{i}$ from the first stage, and semantic BEV features $\{F_{\text{s}_{i}}^{\text{BEV}}\}_{i}$ from the second stage.
The BEV encoder downsamples $F^{\mathrm{BEV}}_{c}$ through convolutional blocks, each followed by a VRWKV block, and integrates geometric BEV features $F_{\text{g}_{i}}^{\text{BEV}}$ and semantic BEV features $F_{\text{s}_{i}}^{\text{BEV}}$ using an ARF module [65]. The process can be represented as Eq. (10):
$F^{\text{BEV}}_{f_{i}}=\smashoperator[]{\sum_{k\in\{\text{g},\text{s},\text{c}\}}^{}}\Big{(}F^{\text{BEV}}_{k_{i}}\odot A^{\text{BEV}}_{k_{i}}\Big{)},$ (10)
where $F^{\text{BEV}}_{k_{i}}$ and $A^{\text{BEV}}_{k_{i}}$ represent the $k$-th class feature map of the $i$-th BEV feature and its corresponding attention weight, respectively. The decoder restores spatial resolution using transposed convolutions and skip connections. The processed BEV features are reshaped and upsampled to generate a voxel-level representation $\hat{Y}_{\text{$b$}}$.
For voxel-level feature processing, we adopt the approach of SGN [66] for voxel handling. Specifically, we fuse and refine three types of features: $F_{\text{s}}^{3D}$ obtained from semantic refinement, non-seed voxel features $F_{\text{n}}^{3D}$ filtered by the proposal network, and $F^{3D}_{o}$ from the proposal network. This procedure is formally expressed in Eq. (11):
$\hat{Y}_{\text{$v$}}=\text{MSSP}([F_{\text{s}}^{3D},F_{\text{n}}^{3D},F_{\text{$o$}}^{3D}]),$ (11)
where MSSP denotes the multi-scale semantic propagation.
Subsequently, $\hat{Y}_{\text{$v$}}$ is concatenated with $\hat{Y}_{\text{$b$}}$ and processed through a 1D convolutional layer to generate the final unified representation $\hat{Y}_{f}$. We optimize the model using a combination of scene- and class-wise semantic loss $L_{sem}^{scal}$, geometry $L_{geo}^{scal}$, and cross-entropy loss $L_{ce}$. The overall loss function is defined in Eq. (12):
$L_{ssc}=L_{sem}^{scal}(\hat{Y},Y)+L_{geo}^{scal}(\hat{Y},Y)+L_{ce}(\hat{Y},Y),$ (12)
where $L_{sem}^{scal}$ ensures semantic consistency at both scene and class levels, $L_{geo}^{scal}$ enforces geometric fidelity, and $L_{ce}$ minimizes classification errors.
Loss Function: The total training loss is defined as follows:
$L=L_{geo}+L_{occ}+L_{sem}+L_{ssc}.$ (13)
4.4 OoD Occupancy Prediction
After achieving high-precision semantic occupancy prediction, we further integrate an out-of-distribution detection mechanism to identify outliers that deviate from the in-distribution regions. For the semantic occupancy prediction results $\hat{Y}_{f}$, we compute the category-specific anomaly score $A(V)$ for each voxel according to the predictive semantics.
Geometry Prior: To more efficiently focus on non-empty regions, we introduce a geometry prior by applying an argmax operation to $\hat{Y}_{f}$. This strategy prioritizes potentially anomalous non-empty regions, significantly improving the effectiveness and accuracy of OoD detection.
VAA-KITTI VAA-KITTI-360
[…]
0.8m 1.0m 1.2m 0.8m 1.0m 1.2m
SGN-S [66] 16.00 24.08 35.09 65.99 17.42 19.99 23.51 66.31
OccOoD-S (ours) 19.64 27.75 38.69 65.91 22.89 25.59 29.21 67.34
Table 1: Quantitative results across different region ranges on the VAA-KITTI and VAA-KITTI-360 datasets. The best results are in bold.
Method IoU
  road
  sidewalk
  parking
  other-grnd.
  building
  car
  truck
  bicycle
  motorcycle
  other-veh.
  vegetation
  trunk
  terrain
  person
  bicyclist
  motorcyclist
  fence
  pole
  traf.-sign
mIoU
SemanticKITTI
VoxFormer [51] 43.21 54.10 26.90 25.10 7.30 23.50 21.70 3.60 1.90 1.60 4.10 24.40 8.10 24.20 1.60 1.10 0.00 13.10 6.60 5.70 13.41
SGN [66] 45.42 60.40 31.40 28.90 8.70 28.40 25.40 4.50 0.90 1.60 3.70 27.40 12.60 28.40 0.50 0.30 0.10 18.10 10.00 8.30 15.76
Symphonies [41] 42.19 58.40 29.30 26.90 11.70 24.70 23.60 3.20 3.60 2.60 5.60 24.20 10.00 23.10 3.20 1.90 2.00 16.10 7.70 8.00 15.04
CGFormer [110] 44.41 64.30 34.20 34.10 12.10 25.80 26.10 4.30 3.70 1.30 2.70 24.50 11.20 29.30 1.70 3.60 0.40 18.70 8.70 9.30 16.63
OccOoD (Ours) 45.11 61.80 31.60 30.50 10.00 27.90 26.60 2.60 1.00 0.80 5.60 26.70 11.90 28.90 0.50 0.80 0.00 18.70 9.40 5.70 15.85
VAA-KITTI
SGN-S [66] 39.07 46.62 26.54 15.88 0.12 18.34 20.51 0.62 0.44 0.35 1.23 24.76 6.40 21.78 0.48 1.29 0.00 12.25 8.70 4.57 11.10
OccOoD-S (Ours) 40.04 50.17 27.42 18.4 0.01 17.38 20.19 0.74 0.26 0.56 1.04 25.13 5.53 23.56 0.67 1.95 0.00 11.35 8.15 5.80 11.49
Table 2: Quantitative results on the SemanticKITTI test set [3] and VAA-KITTI. The best and the second best results are in bold and underlined, respectively.
Semantic-aware Anomaly Scoring: We categorize voxels into two groups based on their semantic predictions: instance (e.g., car, person, bicycle), and region (e.g., road, sidewalk, vegetation). Anomaly scores are computed using methods tailored to each group’s unique characteristics.
  • For instance, we employ cosine similarity to measure deviations in feature space, as these classes often exhibit distinct geometric and semantic features. The cosine similarity for per-voxel’s logit is computed as:
    $S_{\text{cosine}}(v_{i})=1-\cos(\mathbf{l}(v_{i}),\bar{\mathbf{l}}),$ (14)
    where $\mathbf{l}(v_{i})$ is the logit vector of voxel $v_{i}$, and $\bar{\mathbf{l}}$ is the mean logit vector of the same class.
  • For region, we use entropy-based scoring to capture prediction uncertainty, as these classes typically cover larger, more homogeneous areas. The entropy-based anomaly score is computed as:
    $S_{\text{entropy}}(v_{i})=-\sum_{k\in\mathcal{K}}p(k|v_{i})\log p(k|v_{i}),$ (15)
    where $\mathcal{K}$ represents the set of possible classes, and $p(k|v_{i})$ denotes the probability that voxel $v_{i}$ is predicted to be of class $k$. To prioritize instance classes, the anomaly scores for region classes are weighted accordingly. Higher entropy values indicate greater uncertainty, which is often associated with OoD regions.
Ultimately, the anomaly scores $A(V)$ are normalized to $[0,1]$ to ensure comparability across classes. Voxels predicted as empty are assigned the minimum anomaly score, effectively filtering them from further analysis.
5 Experiments
Two sets of experiments are conducted to evaluate the effectiveness of the proposed OccOoD framework. First, we assessed its performance on OoD on established datasets. Then, we evaluated the performance of semantic occupancy prediction on public benchmarks as well as established datasets. After that, we performed several ablation studies to validate the effectiveness of our design. For additional details on datasets, evaluation metrics, and implementation specifics, please refer to the supplementary materials.
5.1 Results on Different Datasets
Figure 4: Visualization of the results on the VAA-KITTI. From left to right are the input RGB images, the SGN-S [66] results, the OccOoD-S (ours) results, the OoD ground truths, and the semantic occupancy prediction ground truths. Zoom in for a better view.
Quantitative Results: In Table 1, the OoD detection results are reported. Across the three different scales of AuPRCr, our network achieves a significant improvement over the baseline. On the VAA-KITTI dataset, our performance improves by more than $3\%$ in terms of AuPRCr. Similarly, on the VAA-KITTI-360 dataset, our method demonstrates a performance gain of over $5\%$. Furthermore, we present the experimental results of semantic occupancy prediction in Table 2. Our method consistently demonstrates performance improvements over the baseline model, regardless of whether the evaluation is conducted on public datasets or our self-constructed datasets. These results fully validate the effectiveness of our designed VBPF, which progressively integrates features from the BEV perspective, thereby enhancing global context perception. For more details on experimental results, please refer to the supplementary materials.
Qualitative Results: In Fig. 4, four different sets of visualization results are presented. It can be observed that without incorporating a top-down perspective, the baseline network exhibits significant estimation biases. Specifically, by integrating two different observation perspectives, our network achieves more accurate localization for small objects (skunk). Meanwhile, for elongated objects (cow), accurately determining their size from a single image alone is inherently challenging. However, by leveraging multi-view observations, the proposed network makes predictions that are more physically consistent and geometrically plausible.
[…]
Table 3: Ablation studies of OccOoD on established datasets.
[…]
Table 4: Ablation and comparison studies of geometry prior on established datasets and regions within 1.2m.
5.2 Ablation Studies
Semantic-aware Anomaly Scoring: Table 3 demonstrates the effectiveness of our proposed Semantic-aware Anomaly Scoring (ASS). Compared to the other three scoring methods, our approach improves AuPRCr by more than $10\%$ across all scales. This is because we incorporate the spatial occupancy characteristics of different objects in 3D space, applying a cosine similarity method for instance-level objects while using an entropy-based approach for larger and more uniform regions. Moreover, in terms of AuROC, we also achieved impressive improvements.
Geometry Prior: In Table 4, we analyze the importance of Geometry Prior by examining AuPRCr and AuROC. With the geometry prior, the network achieves a significant improvement in AuPRCr. In the absence of the geometry prior, the higher AuROC score is due to the fact that it also considers the uncertainty of empty voxel as an anomaly voxel. In our dataset, positive samples are extremely rare, making the AuPRCr metric more suitable for evaluating the model’s performance.
Sequences Used for Training Depth Estimation Model Used for Training Methods #Sequences IoU mIoU Methods Depth Estimation Model IoU mIoU SGN-T [66] 10 46.21 15.32 SGN-S [66] MobileStereoNet [76] 42.76 14.06 OccOoD-T 46.83 16.80 OccOoD-S 43.33 14.54 SGN-T [66] 7 45.15 15.23 SGN-S [66] SPI [45] 41.65 13.68 OccOoD-T 45.49 15.86 OccOoD-S 42.75 13.79
Table 5: Ablation and comparison studies on SemanticKITTI val set [3]. The left table examines different training sequences, whereas the right table studies different depth estimation models.
Impact of Depth Prior: Table 5 demonstrates the impact of different depth models on detection performance. Using only monocular depth estimation, our method outperforms the baseline by $1.1\%$. With the more accurate MSNet3D model [76], the improvement reaches $0.57\%$. The top-down observation perspective mitigates depth map limitations, enabling our network with monocular depth to nearly match the baseline’s performance using stereo-depth maps.
6 Conclusion
In this study, we look into out-of-distribution semantic occupancy prediction to address OoD voxel detection in 3D space. To address the scarcity of anomaly annotations in existing datasets, we developed a Synthetic Anomaly Integration Pipeline and introduced two datasets, VAA-KITTI and VAA-KITTI-360, for evaluating OoD detection in 3D occupancy tasks. We put forward OccOoD to integrate OoD detection into 3D semantic occupancy prediction. By combining voxel-based and BEV representations, OccOoD effectively captures both local details and global context, thereby enhancing its semantic and geometric understanding. Experiments show that OccOoD delivers state-of-the-art OoD detection performance on VAA-KITTI and VAA-KITTI-360, while also maintaining competitive occupancy prediction performance on public benchmarks.
Limitation: The performance of current semantic occupancy prediction methods is limited by geometric inaccuracies in estimated occupancy grids, which adversely affect precision of anomaly detection. We anticipate that future advancements in the field will substantially address these limitations, leading to more robust 3D perception.
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (No. 62473139) and in part by the Open Research Project of the State Key Laboratory of Industrial Control Technology, China (Grant No. ICT2025B20).